OpenStudy (thatcreepyboy):

Medal and Fan. Can someone check my answer? Find the diameter of a cone that has a volume of 83.74 cubic inches and a height of 5 inches. Use 3.14 for pi. 3 inches 4 inches 8 inches 16 inches

OpenStudy (mhchen):

Do you know the formula for the volume of a cone?

OpenStudy (thatcreepyboy):

A cone?

OpenStudy (mhchen):

Yes, if you google it, it's the base * height * (1/3) So the base is a circle, which is r^2 * pi So you have (r^2*pi)*(5) = 83.74 Once you solved for r, the diameter is just radius * 2.

OpenStudy (thatcreepyboy):

I thought it was 8.

Directrix (directrix):

>>I thought it was 8. V = (1/3) * pi * r^2 * h 83.74 = 1/3 * 3.14 *r^2 * 5 r = 4.0016 --> r = 4 to the nearest integer Diameter = 2* 4 = 8 ThatCreepyBoy I agree that 8 is the correct diameter. <-----

OpenStudy (thatcreepyboy):

Oh okay thanks

Directrix (directrix):

You are welcome.